Luke Barrett, Matt Ginn and Jake Youso – Water Runoff at Holy Cross

Your presentation to the class was mostly good, although it certainly did not go as deeply into most of the aspects of this topic as it could have.

The paper is also generally OK, although it is clearly three separate sections spliced together rather than a single text. There are a lot of generalities repeated several times in the different sections. There is less content here than I hoped you would be able to generate.

You also did not include the discussion of how you would take the data from a “finergrained” survey of the campus and use that to identify the problem areas via a statistical analysis. If you look back at my email to the run-off groups from May 3, you will see that I asked everyone to do that. So this is an important omission.

There are also some technical issues. The major thing there is that you have not included in the paper a complete bibliography of the sources you consulted. I hope that you didn’t misunderstand the terms of the assignment. Even though I asked you to do the annotated bibliography earlier, you were still supposed to include a complete bibliography in the final paper since, as was true for several groups, additional sources might have emerged as work on the project proceeded. Plus, citations in the text should be checkable from the paper itself (without referring to what you submitted earlier). I can guess how that probably happened, but that just is not acceptable for a research paper like this. (Be aware, too, that in a lot of courses here at HC, problems with citations would carry extremely steep penalties on the overall grade for the paper.)

Specific Comments

1) Actually, if you look at the history of the environmental movement in the US and other countries, you will see that concern over water and air pollution developed a long time before issues about greenhouse gases and global warming. If you have the chance, look up Rachel Carson’s book Silent Spring.

2) As we discussed several times, it is actually now illegal to allow that kind of waste to flow into nearby rivers. The College is required to dispose of all chemical wastes in approved ways.

3) This is somewhat contradicted by what you say later about the initiative to reduce the amount of salt that is used. This could be a coordination issue between the different sections (authors), or it could be a case of not explaining something that well. I cannot really tell which it is.

4) Where was this photo taken? Was it on campus? Say more!

5) This is presumably a group effort. Say “we” rather than “I” when discussing how the data was analyzed.

6) The point that the College is on a hillside has been made many times before the Preventative Techniques section. It is overkill to say this again here.

7) Storm drains actually usually carry runoff water into nearby streams, etc. So you could argue that the drains are just moving the problem from one location to another.

Final Project Grades Presentation: B (85) Paper Content: B- (82) Paper Form: B (85)